OpenStudy (anonymous):

Solve for the variable. Reduce the answer to lowest terms. -8p = -176

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

p=22

OpenStudy (anonymous):

a answer

OpenStudy (anonymous):

-8p= -176 p = -176/-8 p = 22

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

in maths we have operations which are opposite of each other addition/subtraction multiplication/division REMEMBER WE HAVE TO MAINTAIN EQUALITY, SO WHATEVER ACTION WE TAKE ON ONE SIDE, WE HAVE TO DO THE SAME ON OTHER SIDE TOO!! If u hv to solve 5x +7 = 11 solving means that u hv to separate the variable x frm everything and get its value this means u hv to separate it from +7 and also the 5 of 5x subtraction is opposite of addition so we subtract 7 from both sides 5x+7-7 = 11-7 5X=4 now 5 is multiplying x and opposite of multiply is divide dividing both sides by 5 5x/5 = 4/5 gives x = 4/5

OpenStudy (anonymous):

once u understand the concept we do it in short 5x+7=11 we subtract 7 from both sides and for the left hand side do it mentally and show for right hand side only 5x=11-7 5x=4 then divide both sides by 5 and for the left hand side do it mentally and show for right hand side only x=4/5 got it????? If u r happy with the explanation, pls click the GOOD ANSWER button for me :))